An automotive researcher wanted to estimate the difference in distance required to come to a complete stop while traveling 40 miles per hour on wet versus dry pavement. Because car type plays a role, the researcher used eight different cars with the same driver and tires. The braking distance (in feet) on both wet and dry pavement is shown below. Test whether there is a difference in stopping distance for wet vs. dry pavement. (Paired Sample)

Car

1

2

3

4

5

6

7

8

Wet

106.9

100.9

108.8

111.8

105.0

105.6

110.6

107.9

Dry

71.8

68.8

74.1

73.4

75.9

75.2

75.7

81.0
Answer #1

Ans:

Car 1 2 3 4 5 6 7 8
Wet 106.9 100.9 108.8 111.8 105 105.6 110.6 107.9
Dry 71.8 68.8 74.1 73.4 75.9 75.2 75.7 81
d 35.1 32.1 34.7 38.4 29.1 30.4 34.9 26.9
d-bar 32.7
sd 3.767

Test statistic:

t=(32.7-0)/(3.767/SQRT(8))

t=24.55

df=8-1=7

p-value=0.0000

Reject the null hypothesis and conclude that  there is a significant difference in stopping distance for wet vs. dry pavement.